Aunt Ellen is pricing cakes for a baby shower she is throwing. She wants one large cake shaped like a duckling and also some cupcakes in pastel colors. Georgetown Bakery charges $1 for each cupcake, plus $54 for the large cake. Kayla's Sweet Shoppe charges $27 for the large cake and $4 for each cupcake. If Aunt Ellen orders a certain number of cupcakes, the cost will be the same at either bakery. What would the total cost be? How many cupcakes would that include?


1
Expert's answer
2022-04-22T02:58:37-0400

Total cost in Georgetown Bakery is 54 +x, where x - number of cupcakes.

Total cost in Kayla's Sweet Shoppe is 27 + 4x.

If cost is the same in either factory, then:

54 + x = 27 + 4x,

3x = 27,

x = 9.

So, number of cupcakes is 9.

Total cost is 54 + 9 = $63.